OpenStudy (anonymous):

Ingrid is making a quilt using squares that measure 5 in. on one side. What is the length of of a diagonal of one of the quilt squares. Round to the nearest tenth.

OpenStudy (radar):



OpenStudy (radar):

Use the pythgoreas. \[\sqrt{5^{2}+5^{2}}=\sqrt{50}=5\sqrt{2}=\]Use calculator.

OpenStudy (radar):

Answer will be larger than 5 and less than 10, let me know what you get.

OpenStudy (anonymous):

7?

OpenStudy (anonymous):

I have for answers 15.8: 250: 3.2: and 20!

OpenStudy (radar):

Yes 7.07

OpenStudy (anonymous):

Thats not one of my choices though!

OpenStudy (radar):

You don't have an option 7?

OpenStudy (anonymous):

no i have 15.8; 250; 3.2; and 20.

OpenStudy (radar):

That is the answer using the supplied information. Double check everything? 5" square, problem number etc. Something is amiss.

OpenStudy (radar):

It is requesting the diagonal of one square right??

OpenStudy (anonymous):

yes of one diagonal

OpenStudy (radar):

Problem number matches up with answer number.

OpenStudy (anonymous):

Yes. It does.

OpenStudy (radar):

I would close this question. and then repost it, maybe someone else can figure out what this is all about.

OpenStudy (anonymous):

okay.